OpenStudy (anonymous):

x^2-2x-3 =(x-2)(x-1)? is this correct?

OpenStudy (anonymous):

(x-2)(x-1) =x^2 -3x +2 so its wrong

OpenStudy (anonymous):

thats righttttttttttttttttt!!!!!!!!!!!! therefore its PRIME! DUH!!!!

OpenStudy (anonymous):

The end numbers will add to the coefficient of the middle term and multiply to the last term

OpenStudy (anonymous):

I was adding the 2 and 1 with 3 instead of multiplying it

OpenStudy (mertsj):

-3x1=-3 -3+1=-2

OpenStudy (anonymous):

x^2-2x-3 = (x-3)(x+1)

OpenStudy (anonymous):

I was doing it backwards.

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

the factors of 3 are 1 and 3. The middle term is negative so the bigger factor is negative and the end term is negative so they have to have opposite signs. That's how I always thought about it at least

OpenStudy (anonymous):

so its prime, because it cant be factored. am i right?

OpenStudy (anonymous):

no wait... it can be! x^2-2x-3 = (x-3)(x+1) ??

OpenStudy (anonymous):

Yeppers :)

OpenStudy (anonymous):

i got x^2-3+2